# A Qubit as a Bridge Between Statistical Mechanics and Quantum Dynamics

### v1

@@ Abstract
This work presents a unified perspective on thermal equilibrium and quantum dynamics by examining the simplest quantum system, a qubit, as a foundational model. We show that both the thermal partition function and the Loschmidt amplitude can be understood as extensions of a single analytic function along different paths in the complex plane. The zeros of Loschmidt amplitude encode dynamical features such as orthogonality, rate function singularities, and quantum speed limits, in analogy with the role of partition function zeros in equilibrium statistical mechanics. We further establish, through the Cauchy-Riemann equations, that the high-temperature specific heat corresponds to early-time evolution. The discussion follows a pedagogical progression from a single qubit to an interacting spin chain.
@@ I Introduction
To develop this connection, we begin with a single qubit. We observe that both the partition function and the Loschmidt amplitude can be expressed as a polynomial in a complex variable $y$, whose roots encode unifying features of thermal and dynamical behavior. While the model lacks true phase transitions, the zeros still reflect important signatures such as bounds imposed by quantum speed limits and logarithmic divergences in the rate functions. The framework is then generalised to many-body systems, including a chain of non-interacting qubits and an interacting spin system. Throughout, we take a pedagogical approach, highlighting how basic models can illuminate the analytic structure underlying both equilibrium and nonequilibrium quantum physics.
@@ I.1 Outline
This paper is organized as follows. Section II introduces the central model and describes the physical setup. While Section III discusses the thermal description of the qubit using the canonical partition function, Section IV turns to the quantum dynamics of the same system at zero temperature. We compute the Loschmidt amplitude, and identify special times when the system becomes orthogonal to its initial state. Section V develops an analytic framework that brings together thermal and quantum descriptions and Section VI extends this framework to many-body settings. Starting with a chain of $N$ non-interacting qubits, we move towards interactions by considering an interacting spin chain with open boundary conditions. Section VII concludes with reflections on the significance of the unifying analytic framework as well as its pedagogical potential. Appendix A discusses Mandelstam–Tamm and Margolus–Levitin bounds for our specific models.
@@ II The Two-Level System: Setup
Let us consider a qubit with two non-degenerate energy levels. The Hamiltonian lives in a two-dimensional Hilbert space, with the energy spectrum given by
@@
where $|0\rangle$ and $|1\rangle$ form an orthonormal basis of energy eigenstates, see Fig. 1. This defines the simplest non-trivial quantum system. In quantum information, this system naturally serves as a qubit, the basic unit of quantum computation. Another physical realization is a spin-$\frac{1}{2}$ particle in a magnetic field, where the energy levels correspond to spin-up and spin-down along the field axis.
@@ III Thermal Description: Quantum Statistical Mechanics
The thermodynamic properties of the qubit, coupled to a heat bath at inverse temperature $\beta=1/(k_{B}T)$, where $T$ is the temperature and $k_{B}$ is the Boltzmann constant, are governed by the partition function
$Z(y)=\frac{1}{2}\sum_{\alpha=0}^{1}e^{-\beta E_{\alpha}}=\frac{1}{2}(1+y),\quad y=e^{\beta J},$ (2)
with $E_{0}=-J$ and $E_{1}=0$ as defined in Eq. (1). The prefactor $\frac{1}{2}$ helps in subtracting the entropy of the high-temperature limit. For positive temperatures, the physically relevant regime in this work, the parameter $y$ lies in the range $[1,\infty)$.
Upon including the negative temperatures, a regime associated with population inversion in certain systems [2], the domain of $y$ extends to the interval $(0,\infty)$.
Importantly, for all real values of $\beta$, the function $Z(y)$ remains analytic and strictly positive.
@@ IV Quantum Description: Unitary Time Evolution
$|\psi(0)\rangle=\frac{1}{\sqrt{2}}(|0\rangle+|1\rangle),$ (3)
@@
Here $y=e^{iJt}$ is the natural analytic continuation of the variable defined in Eq. (2). This complex Boltzmann factor traverses a unit circle as $t\in(-\infty,\infty)$ [3] in the complex-$y$ plane, see Fig. 2. The unitary evolution introduces a relative phase between the energy eigenstates, though the individual state probabilities remain constant,
@@
$L(y)=\langle\psi(0)|\psi(t)\rangle=\frac{1}{2}\sum_{\alpha}e^{iE_{\alpha}t}=\frac{1}{2}(1+y).$ (6)
This overlap determines how distinguishable the time-evolved state is from the initial one, when $L(y)=0$, the states are orthogonal and perfectly distinguishable.
@@
$P(y)=|L(y)|^{2}.$ (7)
This quantity oscillates in time, and the qubit returns to its initial state periodically, see Inset B, Fig. 3. These oscillations are analogous to Rabi cycles seen in two-level systems under resonant driving, though here they arise purely from free evolution. The return probability vanishes when $\cos(Jt)=-1$, at
$t_{c}=\frac{(2n+1)\pi\hbar}{J},\quad n\in\mathbb{Z}.$ (8)
@@
$t_{\perp}=\frac{\pi\hbar}{J},\quad(n=0).$ (9)
@@ V A Unified Framework: Partition Function and Loschmidt Amplitude
Figure 2: The blue line represents the real Boltzmann factor from statistical mechanics, which never intersects the isolated zero at $y=-1$ (the green point). The red unit circle corresponds to quantum time evolution. It traces the imaginary Boltzmann factor, periodically crossing the zero. The brown square at $y=1$ marks both the infinite-temperature limit $\beta\to 0$ and the initial time of the quantum system $t\to 0$.
@@ V.1 High-Temperature Limit and Early-Time Expansion: Cauchy-Riemann Equations
We now analyze the behavior of $\mathcal{L}(y)$ in the vicinity of $y=1$, which corresponds to the infinite-temperature limit in the thermal case and the initial time in the quantum evolution. In the thermal regime, the second derivative of the free energy determines the specific heat,
$C(\beta)=-k_{B}\,\beta^{2}\frac{d^{2}f_{\text{thermal}}}{d\beta^{2}}.$ (15)
This gives specific heat at high-temperatures as,
$C(\beta)\,\,\,\,\,\underset{\mathclap{\text{high }T}}{\approx}\,\,\,\,-\frac{k_{B}}{8}\,J^{2}\beta^{2}.$ (16)
Since $f_{\text{dyn}}(y)$ is analytic in a neighborhood of $y=1$, see Eq. (10) and Eq. (12), the Cauchy-Riemann conditions apply and allow us to expand the dynamical free energy $f_{\text{dyn}}(y)$ in a Taylor series around $y=1$. These conditions guarantee that the complex derivatives are independent of the directions along which the point is approached. We choose two orthogonal directions: (i) along the real axis and (ii) along the tangent to the unit circle $y=1$ These correspond, respectively, to high-temperature expansion in $\beta$ and early-time expansion in $t$, see Fig. 2. As a result, just as the high-temperature specific heat reflects energy fluctuations, the same information appears in the coefficient of the $t^{2}$ term in the Taylor expansion of $f_{L}(y)$. This analytic equivalence illustrates how early-time evolution in the quantum regime mirrors high-temperature thermodynamic behavior, see Table 1. In particular, for early times i.e. $\frac{Jt}{\hbar}\ll 1$, this correspondence gives us
$f_{L}(t)\approx C(\beta)\big{|}_{\beta\to it/\hbar}\,\,\,\,\,\underset{\mathclap{\text{small }t}}{\approx}\,\,\,\,\,\,\frac{1}{8\hbar^{2}}J^{2}t^{2},$ (17)
where the second approximation follows from Eq.(15) and Eq.(16). The quadratic behavior in $t$ near $t=0$ is shown as Inset A, Fig. 3.
@@
Variable $y=e^{\beta J}\in\mathbb{R}^{+}$ $y=e^{iJt},\ |y|=1$
@@
Energy fluctuation near $y=1$ $-\beta^{2}\dfrac{d^{2}f_{\text{thermal}}}{d\beta^{2}}\Big{|}_{y=1}$ $t^{2}\dfrac{d^{2}f_{L}}{dt^{2}}\Big{|}_{y=1}$
@@ V.2 Zeros and Divergences
as seen from Eq. (10). The importance of zeros of the partition function was originally explored in the context of phase transitions by Lee, Yang, and Fisher [2], where the limit points of complex zeros on the real axis signal singularities in the thermodynamic limit. In general, the behavior of a polynomial is governed by the location of its zeros. In the thermal case, where $y=e^{\beta E}\in\mathbb{R}^{+}$, the zero at $y=-1$ lies outside the physical domain and remains inaccessible. In contrast, in the quantum case, whenever the evolution path crosses the zero at $y=-1$, $f_{L}(y)$ diverges logarithmically, as shown in Fig. 2 and Fig.3. This indicates that in unitary dynamics, the system becomes orthogonal to its initial state, causing the return probability to vanish and the real dynamical free energy to diverge.
@@
The possibility of crossing zero depends on the system’s energy specturm. For example, in a degenerate two-level system, such as when the excited state $\ E_{1}=0$ as defined in Eq.(1), is two-fold degenerate; the zero lies off the unit circle, and the return probability remains nonzero at all times. That is, the system never evolves into a state orthogonal to its initial configuration.
@@
Figure 3: Periodic logarithmic divergences are observed in the real dynamical free energy $f_{L}(y)$ as a function of scaled time in (i) a qubit and (ii) TFIM with open boundary conditions. Inset A. Early evolution of a system is shown to have a quadration dependance in time for small $t$, (17) . Inset B. The system periodically “forgets” ($P(t)=0$) and “remembers” ($P(t)=1$) its initial state due to unitary evolution. The probabilities $P_{0}(t)$ and $P_{1}(t)$ remain constant throughout.
@@
A side note. This framework also invites an electrostatic analogy, especially natural in two dimensions. The zero of $\mathcal{L}(y)$ at $y=-1$ can be viewed as the location of a point charge in the complex-$y$ plane, and the dynamical free energy $f_{dyn}(y)$, plays the role of the electrostatic potential generated by this charge. While this analogy is not particularly useful for the qubit system discussed here, it becomes a powerful tool in understanding phase transitions in more complex many-body systems.
@@ VI Extension to Many-Body Systems
|$\ket{\Psi(0)}=\prod_{j=1}^{N}\ket{\psi_{j}(0)},$ (21)
where each |$\ket{\psi_{j}(0)}$ is the initial state of qubit $j$ as defined in Eq.(3). Under unitary evolution with a non-interacting Hamiltonian, the state at time $t$ becomes
|$\ket{\Psi(t)}=\prod_{j=1}^{N}\big{(}e^{-iHt/\hbar}\ket{\psi_{j}(0)}\big{)}.$ (22)
The corresponding Loschmidt amplitude for the many-body system is the product of single-qubit amplitudes,
@@
where $\mathcal{L}(y)$ is the amplitude associated with a single qubit, as defined in Eq. (10) and $\mathcal{L}_{N}(y)$ denotes the Loschmidt amplitude of the full system. Following the single-qubit case, we define a quantum analog of the free energy density per qubit in the thermodynamic limit:
@@
The physical observables are then given by, see Eq.(13) and Eq.(14),
$f_{L}(y)=\text{Re}(f_{\text{dyn}}(y)),$ (25)
and the return probability in the large-$N$ limit becomes
$P(y)=\exp\left[-2Nf_{L}(y)\right].$ (26)
The exponential form of Eq. (26) plays a central role in large deviation theory [6], where the probabilities of rare fluctuations decay exponentially with system size, as $P(t)\sim\exp[-NI(t)]$, where $I(t)$ is known as the rate function, defined in the large-$N$ limit as
$I(t)=\lim_{N\to\infty}-\frac{1}{N}\ln P(t).$ (27)
In our context, we identify $I(t)=f_{L}(y)$, which captures the fact that the probability of observing a trajectory that deviates significantly from typical quantum evolution is exponentially suppressed in $N$.
The behavior of the real dynamical free energy density in this non-interacting case is identical to that of the single-qubit system. By bridging thermal and quantum perspectives, this connection enables techniques developed for computing free energies in thermal systems to be adapted, via analytic continuation, to study dynamical properties of time-evolving quantum systems through their associated rate functions.

### v2

@@ Abstract
This work presents a unified perspective on thermal equilibrium and quantum dynamics by examining the simplest quantum system, a qubit, as a minimal model. We show that both the thermal partition function and the Loschmidt amplitude can be understood as extensions of a single analytic function along different paths in the complex plane. The zeros of Loschmidt amplitude encode dynamical features such as orthogonality, rate function singularities, and quantum speed limits, in analogy with the role of partition function zeros in equilibrium statistical mechanics. We further establish, through the Cauchy-Riemann equations, that the high-temperature specific heat corresponds to early-time evolution. The discussion follows a pedagogical progression from a single qubit to an interacting spin chain, all with finite dimensional Hilbert spaces.
@@ I Introduction
To develop this connection, we begin with a single qubit, chosen as a minimal prototypical system described by a finite-dimensional Hilbert space, where the central concepts can be illustrated without much technical overhead. We observe that both the partition function and the Loschmidt amplitude can be expressed as a polynomial in a complex variable $y$, whose roots encode unifying features of thermal and dynamical behavior. While the model lacks true phase transitions, the zeros still reflect important signatures such as bounds imposed by quantum speed limits and logarithmic divergences in the rate functions. The framework is then generalised to many-body systems, including a chain of non-interacting qubits and an interacting spin system. Throughout, we take a pedagogical approach, highlighting how basic models can illuminate the analytic structure underlying both equilibrium and nonequilibrium quantum physics.
@@ I.1 Outline
This paper is organized as follows. Section II introduces the central model and describes the physical setup. While Section III discusses the thermal description of the qubit using the canonical partition function, Section IV turns to the quantum dynamics of the same system at zero temperature. We compute the Loschmidt amplitude, and identify special times when the system becomes orthogonal to its initial state. Section V develops an analytic framework that brings together thermal and quantum descriptions and Section VI extends this framework to many-body settings. Starting with a chain of $N$ non-interacting qubits, we move towards interactions by considering an interacting spin chain with open boundary conditions. Section VII concludes with reflections on the significance of the unifying analytic framework as well as its pedagogical potential. Appendix A discusses the Mandelstam–Tamm and Margolus–Levitin bounds in the context of our specific models. Appendix B discusses the quantum Zeno paradox that is based on the early time behaviour developed in Sec. V.1. Appendix C outlines a project idea and a set of homework exercises designed to reinforce the concepts developed in the main text.
@@ II The Two-Level System: Setup
Let us consider a qubit with $\Omega=2$ non-degenerate energy levels. The Hamiltonian lives in a two-dimensional Hilbert space, with the energy spectrum given by
@@
where $|0\rangle$ and $|1\rangle$ form an orthonormal basis of energy eigenstates, see Fig. 1. This defines the simplest non-trivial quantum system. In quantum information, this system naturally serves as a qubit, the basic unit of quantum computation. Another physical realization of a qubit is a spin-$\frac{1}{2}$ particle in a magnetic field, where the energy levels correspond to spin-up and spin-down along the field axis.
@@ III Thermal Description: Quantum Statistical Mechanics
The thermodynamic properties of the qubit, with $\Omega=2$ states and coupled to a heat bath at inverse temperature $\beta=1/(k_{B}T)$, where $T$ is the temperature and $k_{B}$ is the Boltzmann constant, are governed by the partition function [2],
$Z(y)=\frac{1}{\Omega}\sum_{\alpha=0}^{1}e^{-\beta E_{\alpha}}=\frac{1}{2}(1+y),\quad y=e^{\beta J},$ (2)
with $E_{0}=-J$ and $E_{1}=0$ as defined in Eq. (1). The prefactor $1/\Omega$ simply removes the contribution of the high-temperature entropy, $k_{B}\ln\Omega$, from the free energy. For the case of interest, $J>0$, the parameter $y$ takes values in $1,\infty) [$2]. A useful feature is that the partition function $Z(y)$ is analytic and strictly positive for all real $\beta$.
@@ IV Quantum Description: Unitary Time Evolution
$|\psi(0)\rangle=\frac{1}{\sqrt{\Omega}}(|0\rangle+|1\rangle),\quad(\Omega=2),$ (3)
@@
Here $y=e^{iJt/\hbar}$ is the natural analytic continuation of the variable defined in Eq. (2). This complex Boltzmann factor traverses a unit circle as $t\in(-\infty,\infty)$ [4] in the complex-$y$ plane, see Fig. 2. The unitary evolution introduces a relative phase between the energy eigenstates, though the individual state probabilities remain constant,
@@
$L(y)=\langle\psi(0)|\psi(t)\rangle=\frac{1}{2}\sum_{\alpha}e^{-iE_{\alpha}t/\hbar}=\frac{1}{2}(1+y).$ (6)
This overlap determines how distinguishable the time-evolved state is from the initial one; when $L(y)=0$, the states are orthogonal and perfectly distinguishable.
@@
$P(y)=|L(y)|^{2}=\frac{1}{2}[1+\cos(Jt)]=\cos^{2}(Jt/2),$ (7)
where we used $y=\cos(Jt)+i\sin(Jt)$.
This quantity oscillates in time, and the qubit returns to its initial state periodically. These oscillations are analogous to Rabi cycles seen in two-level systems under resonant driving, though here they arise purely from free evolution. The return probability vanishes when $\cos(Jt)=-1$, at
$t_{c,n}=\frac{(2n+1)\pi\hbar}{J},\quad n\in\mathbb{Z}.$ (8)
@@
$t_{c}=\frac{\pi\hbar}{J},\quad(n=0).$ (9)
@@ V A Unified Framework: Partition Function and Loschmidt Amplitude
Figure 2: The blue line represents the real Boltzmann factor from statistical mechanics, which never intersects the isolated zero at y=1y=-1 (the green point). The red unit circle corresponds to quantum time evolution. It traces the complex Boltzmann factor, periodically crossing the zero. The brown point at $y=1$ marks both the infinite-temperature limit $\beta\to 0$ and the initial time of the quantum system $t\to 0$.
@@
Variable $y=e^{\beta J}\in\mathbb{R}^{+}$ $y=e^{iJt/\hbar},\ |y|=1$
@@
Energy fluctuation near $y=1$ $\beta^{2}\dfrac{d^{2}f_{\text{thermal}}}{d\beta^{2}}\Big|_{y=1}$ $t^{2}\dfrac{d^{2}f_{L}}{dt^{2}}\Big|_{y=1}$
@@ V.1 High-Temperature Limit and Early-Time Expansion: Cauchy-Riemann Equations
We now analyze the behavior of $\mathcal{L}(y)$ in the vicinity of $y=1$, which corresponds to the infinite-temperature limit in the thermal case and the early-time limit in the quantum evolution. In the thermal regime, the second derivative of the free energy determines the specific heat,
$C(\beta)=-k_{B}\,\beta^{2}\frac{d^{2}f_{\text{thermal}}}{d\beta^{2}}.$ (15)
This gives specific heat at high-temperatures as,
$C(\beta)\,\,\,\,\,\underset{\mathclap{\text{ high }T}}{\approx}\,\,\,\,\frac{k_{B}}{4}\,J^{2}\beta^{2}.$ (16)
The $\beta^{2}$ dependence of specific heat in the high temperature regime is generic for systems with a finite dimensional Hilbert space.
Since $f_{\text{dyn}}(y)$ is analytic in a neighborhood of $y=1$, see Eq. (10) and Eq. (12), the Cauchy-Riemann conditions apply and allow us to expand the dynamical free energy $f_{\text{dyn}}(y)$ in a Taylor series around $y=1$. These conditions guarantee that the complex derivatives are independent of the directions along which the point is approached. We choose two orthogonal directions: (i) along the real axis and (ii) along the tangent to the unit circle $|y|=1$. See Fig. 2. These correspond, respectively, to high-temperature expansion in $\beta$ and early-time expansion in $t$. As a result, just as the high-temperature specific heat reflects energy fluctuations which also appears in the coefficient of the $t^{2}$ term in the Taylor expansion of $f_{L}(y)$.
This correspondence links the high-temperature thermodynamic behavior (under $\beta\leftrightarrow t$) to early-time evolution in the quantum regime, see Table 1. In particular, for early times i.e. $\frac{Jt}{\hbar}\ll 1$, this correspondence gives us
$f_{L}(t)\approx\frac{1}{2}\left.\frac{C(\beta)}{k_{B}}\right|_{\beta\to t/\hbar}\,\,\,\,\,\underset{\mathclap{\text{small }t}}{\approx}\,\,\,\,\,\,\frac{1}{8\hbar^{2}}J^{2}t^{2},$ (17)
where the second approximation follows from Eq.(15) and Eq.(16). What is striking is that a high-temperature calorimetric quantity, namely, the specific heat, maps onto the zero-temperature quantum dynamics of the system.
V.1.1 Digression: Quantum Zeno effect
The early-time quadratic dependence of $f_{L}(t)$ is central to understanding the quantum Zeno effect [7, 8, 9]. From Eq. (17) we see that the probability of leaving the initial state grows only quadratically at short times. Independently, any measurement that checks for survival collapses the evolved state onto the initial state. Repeated alternation of short-time evolution with measurement-induced collapse can substantially inhibit the system’s overall evolution. Appendix B provides further discussion.
@@ V.2 Zeros and Divergences
as seen from Eq. (10). The importance of zeros of the partition function was originally explored in the context of phase transitions by Lee, Yang, and Fisher [2], where the limit points of complex zeros on the real axis signal singularities in the thermodynamic limit. In general, the behavior of a polynomial is governed by the location of its zeros. In the thermal case, where $y=e^{\beta E}\in\mathbb{R}^{+}$, the zero at y=1y=-1 lies outside the physical domain and remains inaccessible. In contrast, in the quantum case, whenever the evolution path crosses the zero at y=1y=-1, $f_{L}(y)$ diverges logarithmically, as shown in Fig. 2 and Fig.3. This indicates that in unitary dynamics, the system becomes orthogonal to its initial state, causing the return probability to vanish and the real dynamical free energy to diverge.
@@
The possibility of crossing zero depends on the system’s energy spectrum. For example, in a degenerate two-level system, such as when the excited state $E_{1}=0$ as defined in Eq.(1), is two-fold degenerate; the zero lies off the unit circle, and the return probability remains nonzero at all times. That is, the system never evolves into a state orthogonal to its initial configuration. (See Appendix C.)
@@
Figure 3: Periodic logarithmic divergences are observed in the real dynamical free energy $f_{L}(y)$ as a function of scaled time in (i) a qubit (pink) and (ii) TFIM with open boundary conditions (green).
@@
A side note. This framework also invites an electrostatic analogy, that is especially natural in two dimensions. The zero of $\mathcal{L}(y)$ at y=1y=-1 can be interpreted as the location of a point charge in the complex-$y$ plane, with the dynamical free energy $f_{dyn}(y)$ playing the role of the electrostatic potential generated by this charge [10, 11]. For the simple qubit system, this analogy is mainly illustrative, but in more complex many-body systems, it becomes a powerful way to understand how phase transitions arise [12]. The qubit example nevertheless conveys the essential idea, that the zeros in the complex $y$-plane act like charges whose positions determine both the statistical mechanical behaviour and the zero-temperature quantum evolution.
V.3 On initial state
The particular initial state of Eq. (3) is motivated by the high temperature limit where the two states occur with equal probability. It is also useful to consider a more general normalized state,
$|\psi(0)\rangle=\frac{1}{\sqrt{|c_{0}|^{2}+|c_{1}|^{2}}}(c_{0}|0\rangle+c_{1}|1\rangle),$ (19)
where $c_{0},c_{1}$ are arbitrary complex amplitudes.
Following the same steps as before, the Loschmidt amplitude takes the form
$L(y)=g_{1}+g_{0}\,y,\quad{\textrm{ with }}g_{0}=\frac{|c_{0}|^{2}}{|c_{0}|^{2}+|c_{1}|^{2}},$ (20)
and $g_{1}=1-g_{0}.$ For real $y\geq 1$ and $0<g_{0}<1$, it is possible to interpret this $L(y)$ as a partition function with nonthermal statistical weights for the states. Note that we recover Eq. (6) for $g_{0}=1/2$, whereas the degenerate two level system considered in the previous subsection corresponds to $g_{0}=1/3$.
The zero of $L(y)$ is
$y_{0}=-\frac{1-g_{0}}{g_{0}},$ (21)
which lies on the unit circle only when $g_{0}=1/2$. Thus, for a generic initial state, the Loschmidt amplitude never vanishes; instead, the return probability exhibits a finite minimum at the time $t=tc$ given by Eqs. (8) and (9). The degree of “non-orthogonalizability” can be quantified by the maximum of $f_{L}(y)$ which occurs at $t=t_{c}$. Denoting the minimum distance of the zero, $y_{0}$, from the unit circle by
$\Delta=|y_{0}+1|=\frac{|2g_{0}-1|}{g_{0}},$
we find that the maximum value $f_{\textrm{m}ax}$ depends on $\Delta$ as
$f_{\textrm{m}ax}\sim-\ln\Delta,{\textrm{ as }}\Delta\to 0.$ (22)
The high-temperature specific heat continues to exhibit a $\beta^{2}$ dependence, with a coefficient that depends on $g_{0}$. Importantly, the correspondence between the specific heat and the early-time behavior of the quantum dynamics remains unchanged.
Even in this more general setting, the thermal behavior may still be viewed through the electrostatic analogy: $f(y)$ corresponds to the two-dimensional potential generated by a point charge located at $y=y_{0}.$
@@ VI Extension to Many-Body Systems
[…]
The corresponding Loschmidt amplitude for the many-body system is the product of single-qubit amplitudes, allowing us to write
@@
where $\mathcal{L}(y)$ is the dynamical partition function associated with a single qubit, as defined in Eq. (10) and $\mathcal{L}_{N}(y)$ denotes the corresponding quantity for the full system. Following the single-qubit case, we define a quantum analog of the free energy density per qubit in the thermodynamic limit:
@@
The return probability in the large-$N$ limit becomes
$P(y)=\exp\left[-2Nf_{L}(y)\right],$ (29)
where $f_{L}(y)$ is given by Eq.(13). The exponential form of Eq. (29) plays a central role in large deviation theory [13], where the probabilities of rare fluctuations decay exponentially with system size.